OpenStudy (anonymous):

WILL MEDAL!! The sum of two numbers is 30.the first number 1/5 of the second number. What are the numbers ?

OpenStudy (anonymous):

let first no.=x then second no.=5x x+5x=30 then find x and 5x

OpenStudy (anonymous):

so solve for x+5x=30 I have to find the smaller number and the larger number of the two numbers

jimthompson5910 (jim_thompson5910):

what did you get when you solved `x+5x=30`

OpenStudy (anonymous):

x=5

jimthompson5910 (jim_thompson5910):

correct, now use x = 5 to compute the value of 5x